OpenStudy (anonymous):

What is the solution to this equation? 7(x + 5) - 14 = 28 A. x = 11 B. x = 6 C. x = 1 D. x = 4

OpenStudy (anonymous):

C. x = 1

OpenStudy (anonymous):

7(x + 5) - 14 = 28 7x + 35 - 14 = 28 7x +35 = 42 7x =7 x=1 BUT i guess you just want the answer.. so.